# Vector-Valued Holomorphic Functions and Abstract Fubini-Type Theorems

### v1

@@ 5. Mattner’s results
Theorem 4.2 and in particular the Fubini-type result in assertion d) may be surprising on first glance, since the joint measurability of the function $f$ is not assumed (and not needed in the proof). However, as the following result shows, joint measurability is automatic.
Lemma 5.1 (Mattner [15, p.33]).
@@
Joint measurability results appear to have a long tradition, see e.g. [8, p.42], [5], [1, Lemma 4.51]. However, they are seldom mentioned in courses on measure theory and not widely known. Mattner’s version appears to be the strongest so far where measurability in one variable is paired with continuity in the other.
@@
Theorem 5.3 (Mattner [15, p.32]).
Let $(\Omega,\Sigma,\mu)$ be a measure space, let $O\subset\mathbb{C}$ be open, and let $f:O\times\Omega\to\mathbb{C}$ be a function subject to the following assumptions:
@@
Assertion [C3] in Theorem 5.3 is covered (literally) by part b) of Theorem 4.2. The joint measurability assertion in [C1] follows directly from a) and Lemma 5.1; and the remaining part of [C1] follows since the supremum is effectively a supremum over a countable subset of $A$. (Mattner employs the same argument). Finally, [C2] is a straightforward consequence of a) and the following general result.
@@
By adapting the proof, Lemma 5.4 can be generalized from Banach lattices to ordered Banach spaces with generating positive cone, but we refrain from doing so here. The result may be known, but we do not know of a direct reference, hence we have included the simple proof. Basically, the argument is present in the proof of [16, Lemma on p.72], already mentioned in Remark 4.4, 2).
In the remainder of this section, we show how Theorem 4.2 can be derived from Theorem 5.3, as long as one supposes in addition that $\varphi$ is integration against a finite measure. Since the latter is no restriction, by Remark 3.2, this yields an alternative proof of Theorem 4.2 for the case $p{=}1$. (With a little more effort, one would also obtain a proof for the case $p\geq 1$.) Although Mattner only treats the single-variable case, we expect his approach to work also for functions of several variables.

### v2

@@ 5. Mattner’s results
Theorem 4.3 and in particular the Fubini-type result in assertion d) may be surprising on first glance, since the joint measurability of the function $f$ is not assumed (and not needed in the proof). However, as the following result shows, joint measurability is automatic, at least in the case we consider here ($O\subseteq\mathbb{C}^{d}$).
Lemma 5.1 (Mattner [18, p.33]).
@@
Joint measurability results appear to have a long tradition, see e.g. [11, p.42], [5], [1, Lemma 4.51]. However, they are seldom mentioned in courses on measure theory and not widely known. Mattner’s version appears to be the strongest so far where measurability in one variable is paired with continuity in the other.
@@
Theorem 5.3 (Mattner [18, p.32]).
Let $(\Omega,\Sigma,\mu)$ be a measure space, let $O\subseteq\mathbb{C}$ be open, and let $f:O\times\Omega\to\mathbb{C}$ be a function subject to the following assumptions:
@@
Assertion 3 in Theorem 5.3 is covered (literally) by part b) of Theorem 4.3. The joint measurability assertion in C1 follows directly from a) and Lemma 5.1; and the remaining part of C1 follows since the supremum is effectively a supremum over a countable subset of $A$. (Mattner employs the same argument). Finally, C2 is a straightforward consequence of a) and the following general result.
@@
By adapting the proof, Lemma 5.4 can be generalized from Banach lattices to ordered Banach spaces with generating positive cone, but we refrain from doing so here. The result may be known, but we do not know of a direct reference, hence we have included the simple proof. Basically, the argument is present in the proof of [20, Lemma on p.72], already mentioned in Remark 4.5, 2).
In the remainder of this section, we show how Theorem 4.3 can be derived from Theorem 5.3, as long as one supposes in addition that $\varphi$ is integration against a finite measure. Since the latter is no restriction, by Remark 3.3, this yields an alternative proof of Theorem 4.3 for the case $p{=}1$. (With a little more effort, one would also obtain a proof for the case $p\geq 1$.) Although Mattner only treated the single-variable case, we expect his approach to work also for functions of several variables.
@@
Acknowledgements: This paper was finalized during a research stay of the second author at the Dipartimento di Matematica “Guiseppe Peano” of the Universitá di Torino, Italy. Markus Haase is grateful to this institution for its hospitality and, in particular, to Jörg Seiler for his kind invitation.
Both authors thank the anonymous referee for his valuable remarks, in particular for his references to the works of Grosse-Erdmann and Mujica, which led to a considerable improvement.